OpenStudy (anonymous):

Transform a formula? Medal for help. Solve for the variable c in this equation: a(c + b) = d

OpenStudy (mathmale):

You want to "isolate" c. First, let's move that factor out of the left side and into the right side. How? Divide both sides of the equation by a. type out your response here.

OpenStudy (anonymous):

a(c+b)/a = d/a right?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

a(c+b)/a = d/a can be simplified to c+b =(d/a). So you're right so far. What is your next step? Why?

OpenStudy (anonymous):

b from both sides

OpenStudy (anonymous):

c=d/a-b?

OpenStudy (mathmale):

Looks great. I'd like your result even better if you were to enclose d/a in parentheses: c=(d/a)-b

OpenStudy (anonymous):

Ok but why do I enclose them in parentheses?

OpenStudy (anonymous):



OpenStudy (anonymous):

?